### v3

@@ Abstract
Alzheimer’s disease (AD) is the most common neurodegeneration, annually diagnosed in millions of patients. The present medicine scenario still finds challenges in the exact diagnosis and classification of AD through neuroimaging data. Traditional CNNs can extract a good amount of low-level information in an image but fail to extract high-level minuscule particles, which is a significant challenge in detecting AD from MRI scans. To overcome this, we propose a novel Granular Feature Integration method to combine information extraction at different scales combined with an efficient information flow, enabling the model to capture both broad and fine-grained features simultaneously. We also propose a Bi-Focal Perspective mechanism to highlight the subtle neurofibrillary tangles and amyloid plaques in the MRI scans, ensuring that critical pathological markers are accurately identified. Our model achieved an F1-Score of 99.31%, precision of 99.24%, and recall of 99.51%. These scores prove that our model is significantly better than the state-of-the-art (SOTA) CNNs in existence.
@@ I Introduction
Characterized by chronic neurodegeneration, Alzheimer’s disease causes systematic brain neuron deterioration. It affects the cognitive abilities as well as reduces the individual’s ability to perform simple tasks, and also turns out to be the biggest cause of dementia in older people [1]. AD makes an individual suffer from memory loss, a problem in speaking or the inability to understand, etc. AD also results in the presence of abnormal masses (amyloid plaques) and knotted strings of fibers known as neurofibrillary tangles of the brain. Besides, disconnection of neurons is observed. These are considered to be the hallmark features of AD [2]. The disease is primarily targeting regions of the brain responsible for cognition, memory, and communication. The hippocampus and the entorhinal cortex of the brain, the parts that create a memory, are also among the earliest to be targeted. As one dies, yet another area of the brain suffers and begins to shrink. Because the symptoms of AD vary so dramatically between people, the system then appears even more severe in its signs. The underlying cause of AD remains still, not well known, and widely it is given that there might be an amalgamation of genetic factors along with lifestyle and environmental conditions.
@@
Age is considered as the primary risk factor, recording a majority of the cases in those above 65 years in age. Genetics play a major role, as a family history of AD also increases the risk. AD is diagnosed with the help of a very exhaustive procedure of evaluating the family history, physical examination, cognitive and neurological tests. MRI scans, along with doctors’ help, are of great importance in determining AD. These MRI scans give an accurate image of the brain by describing their shape and size. Comparing them with the history of the patient, memory decline and other cognitive functions are considered to be the primary symptoms of dementia [3]. In this work, we aim to contribute to the advancing state-of-the-art in detecting AD through a proposed method that is efficient in improving the quality of medical diagnosis by adopting DL algorithms.
Such complex DL algorithms are provided by CNNs that can be used with the correct identification of AD. DL has a major influence on the detection of AD that can be highly capable of the analysis of a wide range of data. But what is actually very important is the ability of DL algorithms to process complex data and images obtained from MRI scans in order to look for patterns of degeneration in the brain. Some options that are effective for finding patterns, such as in images, are indeed CNNs. It is thus quite likely that we will diagnose AD in its early stages with the help of these advanced algorithms, and this may slow down the advance towards dementia [4]. The DL algorithms can automatically scan MRI scans from patients. DL approaches appear to achieve improvement in the form of performance and hold promising results for the task of AD identification [5].
The DL model, BFPCNN, introduced in this research, brings Bi-Focal Perspectives and Granular Feature Integration for robust extraction of high-level and low-level features to help healthcare professionals in fast and efficient identification of abnormalities in the brain. We utilized the open-access series of imaging studies, OASIS-1 dataset [6], which includes cross-sectional MRI data acquired on non-demented young, demented older adults, and middle-aged. Sample images from each class of the dataset is shown in Figure 1. Our research is then compared with that previously conducted and other high-performance CNN models. Consequently, our results show that our model outperforms previous works along with SOTA CNNs, highly accurately extracting features and classifying objects.
@@ III-B Methodology
We first preprocessed the augmented MRI images using four preprocessing methods, namely histogram equalization, median filtering, image resizing, and pixel normalization. Histogram equalization improves the images, thus our model can easily analyze these images. Median filtering reduces the noise within an image but does not blur the edges of an important structure. Resizing the dimension of all images passed to the model ensures that it receives images of equal dimensions. Normalization contributes to fast training of the model by eliminating unacceptably large ranges of input values, which might slow down learning. We then used the above-preprocessed images to train our proposed classifier model and classify the MRI images among one of the four AD types. We then evaluated the classification results using metrics like F1-score, recall, accuracy, and precision. The flow of work has been depicted in Figure 2.
@@ 7. Return $normalised\_image$
III-D Proposed Bi-Focal Perspective Convolutional Neural Network (BFPCNN)
@@ 9. $z\leftarrow Dropout(z)$
This feature map then passes through separable convolutional blocks with multiple residual connections [25] ensuring proper gradients flow through the neural network. Separable convolutional layers split the convolutions into depthwise and pointwise parts, which ensure minimal computational complexity while retaining the capability to capture intricate spatial dependencies from the MRI scan. Batch normalization normalizes the activations of the layer, ensuring that the learning is stabilized and the internal covariate shifts are reduced. ReLU activation functions induce a nonlinearity in the models. That must be necessary for extracting more complex relationships from the features extracted.
@@
We introduce Granular Feature Integration to enhance the performance of CNN models by combining both residual connections and multi-scale inception blocks. Residual connections result in smooth flow through the network and solve the vanishing gradient problems associated with inefficient learning. On the other hand, the multi-scale features extracted by the inception blocks enable the model to extract data at different levels of detail, a move that would give better possibilities of analyzing complex patterns in data. Granular Feature Integration provides a powerful method for feature extracting, therefore, resulting in greater performance at the task of AD.

### v7

@@ Abstract
Being the most commonly known neurodegeneration, Alzheimer’s Disease (AD) is annually diagnosed in millions of patients. The present medical scenario still finds the exact diagnosis and classification of AD through neuroimaging data as a challenging task. Traditional CNNs can extract a good amount of low-level information in an image while failing to extract high-level minuscule particles, which is a significant challenge in detecting AD from MRI scans. To overcome this, we propose a novel Granular Feature Integration method to combine information extraction at different scales along with an efficient information flow, enabling the model to capture both broad and fine-grained features simultaneously. We also propose a Bi-Focal Perspective mechanism to highlight the subtle neurofibrillary tangles and amyloid plaques in the MRI scans, ensuring that critical pathological markers are accurately identified. Our model achieved an F1-Score of 99.31%, precision of 99.24%, and recall of 99.51%. These scores prove that our model is significantly better than the state-of-the-art (SOTA) CNNs in existence.
@@ I Introduction
Characterized by chronic neurodegeneration, Alzheimer’s disease causes systematic brain neuron deterioration. It affects the cognitive abilities as well as reduces the individual’s ability to perform simple tasks thereby turning out to be the biggest cause of dementia in older people [1]. An individual diagnosed with AD suffers from memory loss, problems in speaking, inability to understand, etc. Apart from these symptoms, AD also results in the formation of abnormal masses (amyloid plaques) and knotted strings of fibers known as neurofibrillary tangles of the brain. Furthermore, disconnection of neurons is observed as well constituting to the hallmark features of AD [2]. The disease primarily targets regions of the brain responsible for cognition, memory, and communication. The hippocampus and the entorhinal cortex of the brain, known for their responsibility to create a memory, are among the earliest to be targeted. The underlying cause of AD remains is still not well known, it is speculated that it might be an amalgamation of genetic factors along with lifestyle and environmental conditions.
@@
Age is considered to be the primary risk factor since majority of positive cases are recorded in patients aged over 65 years. Genetics play a crucial role, since a family history of AD also increases the risk. Diagnosis of AD is a tedious task involving exhaustive procedures like evaluating the family history, physical examination, cognitive and neurological tests. MRI scans with doctors’ assistance are given great importance to diagnose AD. These MRI scans provide an accurate image of the brain including the description of their shape and size. Comparing them with the history of the patient, memory decline and other cognitive functions are considered to be the primary symptoms of dementia [3]. In this work, we aim to contribute to the advancing AD detection by development of a state-of-the-art mechanism through our proposed method, thereby efficiently improving the quality of medical diagnosis by adopting Deep Learning (DL) algorithms.
Complex DL algorithms have a major influence in detecting AD due to their capability to analyse wide range of data. But the underlying reason for utilization of DL algorithms is because of its ability to process and interpret MRI scans to identify patterns of brain degeneration. Convolutional Neural Networks (CNNs), in particular can recognize such intricate patterns in image data which in association with advanced algorithms enable early diagnosis of AD and contribute in assisting with slowing down its progression towards dementia [4]. DL models provide an automatic analysis of MRI scans from patients, improving diagnostic efficiency while yielding promising results in AD identification [5].
The DL model, BFPCNN, introduced in this research article, utilizes Bi-Focal Perspectives and integrates Granular Features for robust extraction of high-level and low-level features to help healthcare professionals to assist healthcare professionals to identify abnormalities in the brain in an efficient manner while consuming lesser duration at the same time. We utilized the open-access series of imaging studies, OASIS-1 dataset [6], which includes cross-sectional MRI data acquired from nondemented young, demented older adults, and the middle-aged. Sample images from each class of the dataset is shown in Figure 1. Our research is then compared with that previously conducted and other high-performance CNN models. Consequently, our results show that our model outperforms previous works along with SOTA CNNs, highly accurately extracting features and classifying the MRI scans.
@@ III-B Methodology
We first preprocessed the augmented MRI images using four preprocessing methods, namely histogram equalization, median filtering, image resizing, and pixel normalization. Histogram equalization is employed to improve the contrast of the images. Median filtering reduces the noise within an image but does not blur the edges of an important structure. Resizing the dimension of all images passed to the model ensures that it receives images of equal dimensions. Normalization contributes to fast training of the model by eliminating unacceptably large ranges of input values, which might slow down learning. We then used the above-preprocessed images to train our proposed classifier model and classify the MRI images among one of the four AD types. We then evaluated the classification results using metrics like F1-score, recall, accuracy, and precision. The flow of work has been depicted in Figure 2.
@@ 7. Return $normalised\_image$
III-D Granular Feature Integration
Detecting Alzheimer’s Disease (AD) through MRI scans demands models capable of extracting hierarchical features =$\mathcal{F}=\{f_{1},f_{2},\dots,f_{n}\}$ that span both macro- and micro-level brain structures. Conventional Convolutional Neural Networks (CNNs), formulated as $f_{\theta}(x)=\sigma(W*x+b)$ with convolutional weights $W$ and biases $b$, often excel at low-level feature extraction but struggle with high-level granular details critical for AD detection. To overcome this, we propose Granular Feature Integration (GFI), a mechanism that fuses residual learning $R(x)=x+\mathcal{H}(x)$ with multi-scale processing $M(x)=\bigoplus_{s\in S}\psi_{s}(x)$, where $S$ denotes the set of scales and $\psi_{s}$ represents convolutional transformations at scale $s$.
Formally, GFI can be represented as a composite function:
$GFI(x)=R(I(x)),$ (5)
where $R(x)=x+F(x,W)$ denotes the residual mapping and $I(x)=\bigoplus_{k=1}^{n}\delta_{k}(x)$ aggregates multi-scale features through convolutions $\delta_{k}$ with kernel sizes $k\in\{1,3,5,7\}$. The residual pathways ensure that the gradient $\nabla_{x}L$ remains bounded, thus solving the vanishing gradient problem for deep architectures, while the inception modules allow the network to process spatial hierarchies by varying the receptive field $r_{k}$.
By integrating these two mechanisms, the model captures a more expressive feature space:
$\mathcal{F}_{\text{granular}}=\{f_{\text{global}},f_{\text{local}}\},$ (6)
where $f_{\text{global}}$ represents macroscopic brain structures, and $f_{\text{local}}$ encodes subtle neurodegenerative markers.
III-E Proposed Bi-Focal Perspective Convolutional Neural Network (BFPCNN)
@@ 9. $z\leftarrow Dropout(z)$
This feature map then passes through separable convolutional blocks with multiple residual connections [25] ensuring proper gradients flow through the neural network. Separable convolutional layers split the convolutions into depthwise and pointwise parts, which ensure minimal computational complexity while retaining the capability to capture intricate spatial dependencies from the MRI scan. Batch normalization normalizes the activations of the layer, ensuring that the learning is stabilized and the internal covariate shifts are reduced. ReLU activation functions induce a nonlinearity in the models. This is necessary for extracting more complex relationships from the features extracted.